OpenStudy (anonymous):

Solve the initial value problem. dv/dt + rv = -g, where v(0) = v(o)

OpenStudy (mimi_x3):

linear ode?

OpenStudy (anonymous):

yeah.. gravitational acceleration, and velocity

OpenStudy (anonymous):

yes it is linear and integrating factor swill be used .

OpenStudy (mimi_x3):

\[\frac{dv}{dt} +rv =-g\] integrating factor: \(u=e^{\int r(t)dt}\)

OpenStudy (mimi_x3):

hmm..i dont think that is right :/

OpenStudy (mimi_x3):

It's dv/dt

OpenStudy (anonymous):

yup :P

OpenStudy (anonymous):

saw right now .

OpenStudy (mimi_x3):

I assume it's\[\frac{dv}{dt} +r(t)v=-g\]

OpenStudy (mimi_x3):

It is linear!

OpenStudy (mimi_x3):

I hate it when people delete!! :P

OpenStudy (anonymous):

me too :D

OpenStudy (mimi_x3):

why dv/dt=g??

OpenStudy (anonymous):

its acceleration .

OpenStudy (mimi_x3):

isnt acceleration d^2v/dt^2?? havent done there in a while lol

OpenStudy (mimi_x3):

these*

OpenStudy (anonymous):

no derivative of velocity is acceleration .

OpenStudy (anonymous):

me too doing after long time !

OpenStudy (mimi_x3):

yeah!! lol..i forgot

OpenStudy (mimi_x3):

but why does acceleration =g?

OpenStudy (anonymous):

its i guess related to gravitaional field g is gravitational acceleration .

OpenStudy (mimi_x3):

i thought g denotes gravity :/

OpenStudy (anonymous):

so everyone... thank you so much for sticking with this, i have had some issues going on here at the house... if it helps here is the solution... i am trying to understand how to get there... v(t) =(−g/r)+(v(0) +g/r)e^−rt

OpenStudy (anonymous):

dv/dt + rv = -g, where v(0) = v(o) v'(t)+v(t)*r=-g v'(t)=-g-v(t)r v'(t) ----- =1 -g-v(t)r Integrate both side

OpenStudy (anonymous):

Integral of a'(t)/a(t)= log(a(t))

OpenStudy (mimi_x3):

but why isnt it a linear ode??

OpenStudy (mimi_x3):

but isnt**

OpenStudy (anonymous):

i am unsure how you integrate that... integrating the right side is just v(t) right but what about the left

OpenStudy (anonymous):

ok here we go .

OpenStudy (anonymous):

why did you set h(x) = to the integral of rdt

OpenStudy (anonymous):

to get integrating factor.

OpenStudy (anonymous):

okay, i see, cuz it is infront of the v in the original equation.

OpenStudy (anonymous):

yes exactly.

OpenStudy (anonymous):

i am not seeing how you went from your second step... after you multiplied by the integrating factor, to the third line, where you simplified that left hand side.

OpenStudy (anonymous):

ok by product rule \[\Large \frac{d}{dt}(e^{rt}*v)=e^{rt}*\frac{dv}{dt}+v*\frac{d}{dt}(e^{rt})\] \[\Large \frac{d}{dt}(e^{rt}*v)=e^{rt}*\frac{dv}{dt}+e^{rt}*rv\] which is the left hand side .

OpenStudy (anonymous):

i understand when you do that... seems like you are working backwards though, i would of never thought of that.... but the rest of it I understand good. very clear. the next part of the problem ask for b) show that lim v(t) = -g/r, as t goes to pos infinity..... which equation are we suppose to use for this?

OpenStudy (anonymous):

its the one that you got before v = -g/r + e^-rtC right

OpenStudy (anonymous):

the last one as t goes to inginity exponetial terms gets zero you ar left with only -g/r

OpenStudy (anonymous):

oh you would use the equation after you substituted the initial velocity? the very last one in red, i mean we would get the same answer as using the one before putting in the initial velocity, so does it really matter?

OpenStudy (anonymous):

yes use the red one .its important .

OpenStudy (anonymous):

okay... and the last part of this problem asks c) integrate v to obtain the height y, assuming an initial height y(0) =y(o)

OpenStudy (anonymous):

are we integrating the red equation?

OpenStudy (anonymous):

why would you integrate ????

OpenStudy (anonymous):

i dont know... that is exactly what the problem tells me to do.

OpenStudy (anonymous):

thats it you got the soluion which is in red . and you find that v =-g/r when t tends to infinity . that was your question .

OpenStudy (anonymous):

yes, that was part of the question too, and the last part says to integrate v to obtain height y.

OpenStudy (anonymous):

ok integrate it.

OpenStudy (anonymous):

so if i integrate that first term -g/r, and i am finding y(t) then do will it just be -gt/r?